OpenStudy (rebeccapink6398):

WILL MEDAL!!!!!!!!!!! Nancy walked 516 yards in 6 minutes and then another 368 yards in 4 minutes. What was Nancy's average speed? A. 88.4 yd/min B. 178 yd/min C. 89 yd/min D. 89.4 yd/min

OpenStudy (futureescape):

A

OpenStudy (rebeccapink6398):

HOW? PLEASE TECH NOT ANSWER

OpenStudy (ddcamp):

\[Average\ Speed=\frac{Total\ Distance}{Total\ Time}\]

OpenStudy (rebeccapink6398):

(OOPS sorry caps lock lol)

OpenStudy (jim766):

find the two speeds , add them together and divide by 2

OpenStudy (ddcamp):

\[S=\frac{516yds+368yds}{6\min+4\min}\]

OpenStudy (ddcamp):

@Jim766 Since she was walking at each speed a different amount of time, finding the average of the two speeds is different than the average speed.

OpenStudy (rebeccapink6398):

86?

OpenStudy (jim766):

yes, I see just do as he has shown...add the numerators , then divide by 10

OpenStudy (jim766):

516 + 368 =

OpenStudy (rebeccapink6398):

902

OpenStudy (rebeccapink6398):

like that?

OpenStudy (jim766):

516 + 368 = 884

OpenStudy (jim766):

now take 884 and divide by 10. do you have a calculator?

OpenStudy (rebeccapink6398):

yes

OpenStudy (rebeccapink6398):

88.4

OpenStudy (jim766):

right!

OpenStudy (ddcamp):

Yup!

OpenStudy (jim766):

sorry if i confused you before..